Ms K is 6 years more than 3 times as old as her daughter. 8 years ago she was 2 years less than 9 times as old as her daughter was then. What are thier ages?
Answer by checkley71(8403)

(3d+6)-8=9(d-8)-2 or 3d-2=9d-72-2=3d-9d=-72-2+2 or -6d=-72 d=12 age of the daughter. 3*12+6=35=6=42 age of Ms. K.
proof
3*12+6-8=9(12-8)-2
36+6-8=9*4-2
42-8=36-2
34=34.